OpenStudy (anonymous):

what is the solution of y-4x=0 and 3x+6y=9?

OpenStudy (anonymous):

im thinking (0,0)

OpenStudy (anonymous):

That i dont think it right because then the secoun equation would look like this 0=9

OpenStudy (anonymous):

Hey can you help with this one @Chineseboy15

OpenStudy (anonymous):

How about you @UditKulka

OpenStudy (anonymous):

y=4x from the first equation right?

OpenStudy (anonymous):

Substitute it in the second equation

OpenStudy (anonymous):

haha i was totally off my bad

OpenStudy (anonymous):

3x + 6(4x)=9 27x=9 x=9/27 x=1/3

OpenStudy (anonymous):



OpenStudy (anonymous):